OpenStudy (anonymous):

simplify the square root of x squared + 8x + 16

OpenStudy (anonymous):

You mean like this?\[\sqrt{x^2+8x+16}\]

OpenStudy (akashdeepdeb):

\(x^2+8x+16\) can be written in the form of a perfect square by using these identities: (a+b)^2 = a^2 + 2ab + b^2 (a-b)^2 = a^2 - 2ab + b^2 Try to factorize your expression, and you'll get the square root of it. Getting this? :)

OpenStudy (anonymous):

@halorazer yeah